# Variations on quantum de Finetti theorems and operator valued Martin boundaries: a Choquet–Deny approach

### v2

@@ 4. Matrix-valued Biane–Choquet–Deny theorem
We fix a compact group $G$. Let $C(G)$ denote the *-algebra of continuous functions on $G$. We denote by $\mathbb{C}[G]\subset C(G)$ the linear span of all matrix coefficients of finite-dimensional unitary representations. Here, by matrix coefficients, we mean continuous functions on $G$ given as $g\in G\mapsto\langle\pi(g)\xi,\eta\rangle$ for any (finite-dimensional) unitary representation $(\pi,\mathcal{H})$ of $G$ and two vectors $\xi,\eta\in\mathcal{H}$. By the definitions of tensor product and contragredient representations, $\mathbb{C}[G]$ becomes a *-subalgebra of $C(G)$. Moreover, $\mathbb{C}[G]$ is a Hopf *-algebra whose comultiplication $\Delta_{G}$, antipode $S_{G}$, and counit $\epsilon_{G}$ are given as follows:
@@
[…]
@@
4.2. Matrix-valued Martin boundaries on the duals of compact groups
First, we mention the perspective of quantum random walks or random walks on non-commutative spaces. See [3] and references therein for more details. From this standpoint, a pair consisting of a unital *-algebra and a state is interpreted as a non-commutative analog of a probability space. Specially, a unital *-algebra serves as an algebra of random variables, while a state represents a linear functional obtained through expectation. Thus, a state is a non-commutative counterpart of a probability measure. Our results in this section concern the non-commutative probability space from $\mathcal{U}(G)$. See also the comment below (4.1).
@@ 4.2. Matrix-valued Martin boundaries on the duals of compact groups
For any finite-dimensional unitary representation $(\pi,\mathcal{K})$ of $G$ and any $\xi\in\mathcal{K}$, let $\omega_{\xi}$ denote the positive linear form on $B(\mathcal{K})$ given by $\omega_{\xi}(x):=\langle x\xi,\xi\rangle$ for any $x\in B(\mathcal{K})$. Then $\omega_{\xi}\circ\pi\in\mathcal{U}(G)_{*,+}$ and moreover any element $\rho\in\mathcal{U}(G)_{*,+}$ has this form. Let $\rho\in\mathcal{U}(G)_{*,+}$ be given. Note first that by Schur orthogonality relations, $\rho$ is finitely supported, i.e., $\rho|_{B(\mathcal{H}_{\alpha})}=0$ holds for any but finitely many $\alpha\in\widehat{G}$. For $\rho|_{B(\mathcal{H}_{\alpha})}$, let $(\sigma_{\alpha},\mathcal{K}_{\alpha})$ be the GNS representation of $B(\mathcal{H}_{\alpha})$ and let $\xi_{\alpha}\in\mathcal{K}_{\alpha}$ such that $\rho|_{B(\mathcal{H}_{\alpha})}(x)=\langle\sigma_{\alpha}(x)\xi_{\alpha},\xi_{\alpha}\rangle$ for any $x\in B(\mathcal{H}_{\alpha})$. Note that $\mathcal{K}_{\alpha}$ is finite-dimensional, and $\mathcal{K}_{\alpha}=\{0\}$ for all but finitely many $\alpha\in\widehat{G}$. Let $\sigma$ be the direct sum
@@
For any $\rho\in\mathcal{U}(G)_{*}$ we define a linear map $P_{\rho}\colon\mathcal{U}(G)\to\mathcal{U}(G)$ by “$P_{\rho}:=(\mathrm{id}\otimes\rho)\circ\Delta_{G}$”. The precise definition is as follows:
@@
$B_{m,\rho}$ is the closed convex hull of $\mathrm{ex}(B_{m,\rho})$.
@@
where $\mathrm{Tr}_{\alpha}$ is the non-normalized trace on $B(\mathcal{H}_{\alpha})$. Then, $\varphi_{c}$ is lower semi-continuos, additive, and positive-homogeneous. Moreover, $C_{\varphi_{c}}:=\{x\in B_{m,\rho}\mid\varphi_{c}(x)\leq 1\}$ is compact since
@@
As in Lemma 3.1, we showed that $B_{m,\rho}$ is a union of its caps in the above proof. Thus, combining with results in [11, Chapter 13], we obtain the following representation theorem:
@@
Every $x\in B_{m,\rho}\backslash\{0\}$ can be represented by a maximal Borel probability measure $M_{x}$ supported on $\mathrm{ex}(B_{m,\rho})$ (in the sense of Proposition 3.1).
We now study the extreme points $\mathrm{ex}(B_{m,\rho})$ of $B_{m,\rho}$ and obtain a matrix-valued extension of a theorem by Biane [1] that is itself a non-commutative version of the Choquet–Deny theorem.
[…]
@@
The set $\mathrm{ex}(B_{m,\rho})$ of extreme points of $B_{m,\rho}$ is contained in $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$. In particular, $B_{m,\rho}$ is the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$.
@@
Let $x\in\mathrm{ex}(B_{m,\rho})$. For any $\nu\in\mathcal{U}(G)_{*,+}$ we have $P_{m,\nu}(x)\geq 0$ by Proposition 4.1. Moreover, $P_{m,\rho}(P_{m,\nu}(x))=P_{m,\nu}(P_{m,\rho}(x))\leq P_{m,\nu}(x)$, and hence $P_{m,\nu}(x)\in B_{m,\rho}$. Since $\rho$ is generating, we have $\nu\leq\sum_{l=0}^{k}a_{l}\rho^{*l}$ for some $a_{l}\geq 0$ ($l=0,\dots,k$) and $k\geq 0$. Thus, we have
@@
Thus, $\hat{\Delta}_{G}(x_{\mu})=x_{\mu}\otimes e=e\otimes x_{\mu}$ holds, and $e$ is group-like. Moreover, there exists $a\in\mathbb{M}_{m}(\mathbb{C})_{+}$ such that $x=a\otimes e$. Since $x$ is $P_{m,\rho}$-subharmonic, $e$ is also $P_{\rho}$-subharmonic, i.e., $e\in\mathcal{E}_{\rho}(G)$. Therefore, we have $\mathrm{ex}(B_{m,\rho})\subset\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$, and hence $B_{m,\rho}$ is contained in the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$. On the other hand, since $\mathcal{E}_{\rho}(G)\subset B_{1,\rho}$ by definition, we have $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)\subset B_{m,\rho}$. Thus, the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$ is contained in $B_{m,\rho}$, that is, they are equal. ∎
@@
In this part, we deepen our understanding of the relation between matrix-valued Martin boundary and the generalized quantum de Finetti theorem. Throughout this section, we fix a finite-dimensional unitary representation $(\Pi,\mathcal{H})$ of $G$ and $\rho\in B(\mathcal{H})_{*,+}$ that is faithful. Let us denote by the same symbol $\rho$ the linear functional on $\mathcal{U}(G)$ defined as $\rho\circ\Pi$. We remark that $\rho$ is generating if $(\Pi,\mathcal{H})$ is also generating $\widehat{G}$, that is, for any $\alpha\in\widehat{G}$ there exists $l\geq 0$ such that $(\pi_{\alpha},\mathcal{H}_{\alpha})$ is a sub-representation of $(\Pi^{\otimes l},\mathcal{H}^{\otimes l})$.
Let us recall that the UHF (uniformly hyperfinite) algebra $B(\mathcal{H})^{\otimes\infty}$ is defined as the $C^{*}$-inductive limit of the $B(\mathcal{H})^{\otimes l}$ ($l\geq 0$). If $\rho$ is a state of $B(\mathcal{H})$, then the infinite tensor product state $\rho^{\otimes\infty}$ on $B(\mathcal{H})^{\otimes\infty}$ is defined. Thus, we obtain the $C^{*}$-probability space $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$. Moreover, we can regard the family of the subalgebras $B(\mathcal{H})^{\otimes l}$ as a filtration of $\sigma$-algebras. Namely, $(B(\mathcal{H})^{\otimes\infty},(B(\mathcal{H})^{\otimes l})_{l=0}^{\infty},\rho^{\otimes\infty})$ is a filtered $C^{*}$-probability space. For every $l\geq 0$ the conditional expectation $E_{l}$ from $B(\mathcal{H})^{\otimes\infty}$ to $B(\mathcal{H})^{\otimes l}$ is also defined as the linear map induced by $\mathrm{id}_{B(\mathcal{H})}^{\otimes l}\otimes\rho^{\otimes k}\colon B(\mathcal{H})^{\otimes l+k}\to B(\mathcal{H})^{\otimes l}$ ($k\geq 0$). Under this point of view, a sequence $(x_{l})_{l=0}^{\infty}$ in $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$ is said to be adapted sub-martingale if $x_{l}\in B(\mathcal{H})^{\otimes l}$ and $E_{l}(x_{l+1})\leq x_{l}$ for every $l\geq 0$. According to (3.1), we can interpret $K_{m,\rho}$ as the set of paths of $\mathbb{M}_{m}(\mathbb{C})$-valued adapted sub-martingale processes. Here, $B(\mathcal{H})$ is naturally identified with $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$. Such an interpretation was given initially by Izumi [6]. In his work, he initiated the non-commutative Poisson boundary theory. In the paper, our setting is slightly different from his work, but we can find a connection between such a non-commutative boundary theory and quantum information theory.
@@ 6. The case $G=U(n)$ and the generalized quantum de Finetti theorem revisited
In what follows, let $(\Pi,\mathbb{C}^{n})$ denote the fundamental representation of $U(n)$ and fix a faithful positive linear functional $\rho$ on $\mathbb{M}_{n}(\mathbb{C})$ ($=B(\mathbb{C}^{n})$). As in the previous section, $\rho$ will also denote the linear functional $\rho\circ\Pi$ on $\mathcal{U}(U(n))$. We also define $\Pi_{m,l}$ and $\Pi_{m,\infty}$ as the same as in the previous section. We remark that even if the fundamental representation is not generating $\widehat{U(n)}$, the following statement holds true (see also Proposition 5.1).
@@
By Proposition 5.1, $\Pi_{m,\infty}(B_{m,\rho})$ is contained in $K_{m,\rho}$. Combining Theorem 3.1 and Lemma 6.1, every element in $\mathrm{ex}(K_{m,\rho})$ can be approximated by $\Pi_{m,\infty}(\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(U(N)))$. Thus, by Lemma 3.1, $K_{m,\rho}$ is contained in $\overline{\Pi_{m,\infty}(B_{m,\rho})}$. ∎

### v3

@@ 4. Matrix-valued Biane–Choquet–Deny theorem
We fix a compact group $G$ and introduce three *-algebras $C(G)$, $\mathbb{C}[G]$, and $\mathcal{U}(G)$. The notations in this section follow the standard text [9] on compact quantum groups. Let $C(G)$ denote the *-algebra of continuous functions on $G$. We denote by $\mathbb{C}[G]\subset C(G)$ the linear span of all matrix coefficients of finite-dimensional unitary representations. Here, by matrix coefficients, we mean continuous functions on $G$ given as $g\in G\mapsto\langle\pi(g)\xi,\eta\rangle$ for any (finite-dimensional) unitary representation $(\pi,\mathcal{H})$ of $G$ and two vectors $\xi,\eta\in\mathcal{H}$. By the definitions of tensor product and contragredient representations, $\mathbb{C}[G]$ becomes a *-subalgebra of $C(G)$. Moreover, $\mathbb{C}[G]$ is a Hopf *-algebra whose comultiplication $\Delta_{G}$, antipode $S_{G}$, and counit $\epsilon_{G}$ are given as follows:
@@
For any $l\geq 1$ let us consider the (algebraic) dual space $\mathcal{U}(G^{l}):=(\mathbb{C}[G]^{\otimes l})^{*}$. We denote by (,)($\cdot\,$$\cdot\,$) the natural dual pairing between $\mathcal{U}(G^{l})$ and $\mathbb{C}[G]^{\otimes l}$. If $l=1$, we simply write $\mathcal{U}(G)$. By the Hopf *-algebra structure of $\mathbb{C}[G]$, the linear space $\mathcal{U}(G)$ also becomes a unital *-algebra. More explicitly, the multiplication and the *-operation are given as
@@
4.2. Matrix-valued Martin boundary on the dual of compact groups
First, we mention the notion of quantum random walks or random walks on non-commutative spaces. See [3] and references therein for more details. From this standpoint, a pair consisting of a unital *-algebra and a state is interpreted as a non-commutative analog of a probability space. More precisely, a unital *-algebra serves as an algebra of random variables, while a state represents a linear functional obtained through expectation. Thus, a state is a non-commutative counterpart of a probability measure. Our results in this section concern the non-commutative probability space from $\mathcal{U}(G)$. See also the comment below (4.1).
@@ 4.2. Matrix-valued Martin boundary on the dual of compact groups
For any finite-dimensional unitary representation $(\pi,\mathcal{K})$ of $G$ and any $\xi\in\mathcal{K}$, let $\omega_{\xi}$ denote the positive linear form on $B(\mathcal{K})$ given by $\omega_{\xi}(x):=\langle x\xi,\xi\rangle$ for any $x\in B(\mathcal{K})$. Then $\omega_{\xi}\circ\pi\in\mathcal{U}(G)_{*,+}$ and moreover any element $\rho\in\mathcal{U}(G)_{*,+}$ has this form. Let $\rho\in\mathcal{U}(G)_{*,+}$ be given. Note first that by Schur orthogonality relations, $\rho$ is finitely supported, i.e., $\rho|_{B(\mathcal{H}_{\alpha})}=0$ holds for all but finitely many $\alpha\in\widehat{G}$. For $\rho|_{B(\mathcal{H}_{\alpha})}$, let $(\sigma_{\alpha},\mathcal{K}_{\alpha})$ be the GNS representation of $B(\mathcal{H}_{\alpha})$ and let $\xi_{\alpha}\in\mathcal{K}_{\alpha}$ such that $\rho|_{B(\mathcal{H}_{\alpha})}(x)=\langle\sigma_{\alpha}(x)\xi_{\alpha},\xi_{\alpha}\rangle$ for any $x\in B(\mathcal{H}_{\alpha})$. Note that $\mathcal{K}_{\alpha}$ is finite-dimensional, and $\mathcal{K}_{\alpha}=\{0\}$ for all but finitely many $\alpha\in\widehat{G}$. Let $\sigma$ be the direct sum
@@
For any $\rho\in\mathcal{U}(G)_{*}$ we define a linear map $P_{\rho}\colon\mathcal{U}(G)\to\mathcal{U}(G)$ by $P_{\rho}:=(\mathrm{id}\otimes\rho)\circ\Delta_{G}$”. The precise definition is as follows:
@@
$B_{m,\rho}$ is the closed convex hull of $\mathrm{exr}(B_{m,\rho})$.
@@
where $\mathrm{Tr}_{\alpha}$ is the non-normalized trace on $B(\mathcal{H}_{\alpha})$. Then, $\varphi_{c}$ is lower semi-continuous, additive, and positive-homogeneous. Moreover, $C_{\varphi_{c}}:=\{x\in B_{m,\rho}\mid\varphi_{c}(x)\leq 1\}$ is compact since
@@
As in Lemma 3.1, we showed that $B_{m,\rho}$ is a union of its caps in the above proof. Thus, by the same proof of Proposition 3.1, we obtain the following representation theorem:
@@
Every $x\in B_{m,\rho}\backslash\{0\}$ can be represented by a Borel probability measure $M_{x}$ on $B_{m,\rho}$ such that $M_{x}$ is supported on any closed subset of $B_{m,\rho}$ which contains $\mathrm{exr}(B_{m,\rho})$.
We now study the extreme points $\mathrm{exr}(B_{m,\rho})$ of $B_{m,\rho}$ and obtain a matrix-valued extension of a theorem by Biane [1] that is itself a non-commutative version of the Choquet–Deny theorem.
Let us introduce the following notions. A non-zero element $e\in\mathcal{U}(G)$ is said to be group-like (or primitive) if $\hat{\Delta}_{G}(e)=e\otimes e$ holds. We denote by $\mathcal{G}(G)$ the subset of all group-like elements of $\mathcal{U}(G)$. We also define by $\mathcal{E}(G):=\mathcal{G}(G)\cap\mathcal{U}(G)_{+}$, the subset of all exponential elements. Then for any $e\in\mathcal{E}(G)$ and $f\in\mathbb{C}[G]$ we have that:
@@
The set $\mathrm{exr}(B_{m,\rho})$ of extreme points of $B_{m,\rho}$ is contained in $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G):=\{a\otimes e\mid a\in\mathbb{M}_{m}(\mathbb{C})_{+},\,e\in\mathcal{E}_{\rho}(G)\}$. In particular, $B_{m,\rho}$ is the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$.
@@
Let $x\in\mathrm{exr}(B_{m,\rho})$. For any $\nu\in\mathcal{U}(G)_{*,+}$ we have $P_{m,\nu}(x)\geq 0$ by Proposition 4.1. Moreover, $P_{m,\rho}(P_{m,\nu}(x))=P_{m,\nu}(P_{m,\rho}(x))\leq P_{m,\nu}(x)$, and hence $P_{m,\nu}(x)\in B_{m,\rho}$. Since $\rho$ is generating, we have $\nu\leq\sum_{l=0}^{k}a_{l}\rho^{*l}$ for some $a_{l}\geq 0$ ($l=0,\dots,k$) and $k\geq 0$. Thus, we have
@@
Thus, $\hat{\Delta}_{G}(x_{\mu})=x_{\mu}\otimes e=e\otimes x_{\mu}$ holds, and $e$ is group-like. Moreover, there exists $a\in\mathbb{M}_{m}(\mathbb{C})_{+}$ such that $x=a\otimes e$. Since $x$ is $P_{m,\rho}$-subharmonic, $e$ is also $P_{\rho}$-subharmonic, i.e., $e\in\mathcal{E}_{\rho}(G)$. Therefore, we have $\mathrm{exr}(B_{m,\rho})\subset\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$, and hence $B_{m,\rho}$ is contained in the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$. On the other hand, since $\mathcal{E}_{\rho}(G)\subset B_{1,\rho}$ by definition, we have $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)\subset B_{m,\rho}$. Thus, the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$ is contained in $B_{m,\rho}$, that is, they are equal. ∎
@@
In this part, we deepen our understanding of the relation between matrix-valued Martin boundary and the generalized quantum de Finetti theorem. Throughout this section, we fix a finite-dimensional unitary representation $(\Pi,\mathcal{H})$ of $G$ and $\rho\in B(\mathcal{H})_{*,+}$ that is faithful. Let us denote by the same symbol $\rho$ the linear functional on $\mathcal{U}(G)$ defined as $\rho\circ\Pi$. We remark that $\rho$ is generating if $(\Pi,\mathcal{H})$ is also generating $\widehat{G}$, that is, for any $\alpha\in\widehat{G}$ there exists $l\geq 0$ such that $(\pi_{\alpha},\mathcal{H}_{\alpha})$ is a sub-representation of $(\Pi^{\otimes l},\mathcal{H}^{\otimes l})$ [^1] For instance, the fundamental representation of $SU(n)$ is generating. On the other hand, it fails for $U(n)$. It is not generating. In fact, the fundamental representation of $U(n)$ does not generate irreducible representations with highest weights containing negative parts. However, it plays the central role in the next section that concerns the relationship separable states of a bipartite quantum system and matrix-valued Martin boundary of random walks on the dual of $U(n)$..
Let us recall that the UHF (uniformly hyperfinite) algebra $B(\mathcal{H})^{\otimes\infty}$ is defined as the $C^{*}$-inductive limit of the $B(\mathcal{H})^{\otimes l}$ ($l\geq 0$). If $\rho$ is a state of $B(\mathcal{H})$, then the infinite tensor product state $\rho^{\otimes\infty}$ on $B(\mathcal{H})^{\otimes\infty}$ is defined. Thus, we obtain the $C^{*}$-probability space $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$. Moreover, we can regard the family of the subalgebras $B(\mathcal{H})^{\otimes l}$ as a non-commutative analogue of filtration of $\sigma$-algebras. Namely, $(B(\mathcal{H})^{\otimes\infty},(B(\mathcal{H})^{\otimes l})_{l=0}^{\infty},\rho^{\otimes\infty})$ is a filtered $C^{*}$-probability space. For every $l\geq 0$ the conditional expectation $E_{l}$ from $B(\mathcal{H})^{\otimes\infty}$ to $B(\mathcal{H})^{\otimes l}$ is also defined as the linear map induced by $\mathrm{id}_{B(\mathcal{H})}^{\otimes l}\otimes\rho^{\otimes k}\colon B(\mathcal{H})^{\otimes l+k}\to B(\mathcal{H})^{\otimes l}$ ($k\geq 0$). Under this point of view, a sequence $(x_{l})_{l=0}^{\infty}$ in $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$ is said to be adapted sub-martingale if $x_{l}\in B(\mathcal{H})^{\otimes l}$ and $E_{l}(x_{l+1})\leq x_{l}$ for every $l\geq 0$. According to (3.1), we can interpret $K_{m,\rho}$ as the set of $\mathbb{M}_{m}(\mathbb{C})$-valued, positive, symmetric, adapted sub-martingale sequences. Here, $B(\mathcal{H})$ is naturally identified with $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$. Such an interpretation was given initially by Izumi [6]. In his work, he initiated the non-commutative Poisson boundary theory. In the paper, our setting is slightly different from his work, but we can find a connection between such a non-commutative boundary theory and quantum information theory.
@@ 6. The case $G=U(n)$ and the generalized quantum de Finetti theorem revisited
In what follows, let $(\Pi,\mathbb{C}^{n})$ denote the fundamental representation of $U(n)$ and fix a faithful positive linear functional $\rho$ on $\mathbb{M}_{n}(\mathbb{C})$ ($B(\mathbb{C}^{n})$). As in the previous section, $\rho$ will also denote the linear functional $\rho\circ\Pi$ on $\mathcal{U}(U(n))$. We also define $\Pi_{m,l}$ and $\Pi_{m,\infty}$ as the same as in the previous section. We remark that even if the fundamental representation is not generating $\widehat{U(n)}$, the following statement holds true (see also Proposition 5.1).
@@
By Proposition 5.1, $\Pi_{m,\infty}(B_{m,\rho})$ is contained in $K_{m,\rho}$. Combining Theorem 3.1 and Lemma 6.1, every element in $\mathrm{exr}(K_{m,\rho})$ can be approximated by $\Pi_{m,\infty}(\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(U(N)))$. Thus, by Lemma 3.1, $K_{m,\rho}$ is contained in $\overline{\Pi_{m,\infty}(B_{m,\rho})}$. ∎